OpenStudy (anonymous):

solve for x x/10= 2 1/5 A. 11/50 B. 22 C. 20 1/5 D. 12 1/5

OpenStudy (anonymous):

B. 22

OpenStudy (anonymous):

thank you so much can you show me how to do it please?

OpenStudy (anonymous):

x/10 = 11/5 you know that 2 1/5 = 11/5 ? do cross-multiplication and you get x = 10 (11/5) = 2*11

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

no worries :)